# Blinded by Generated Contexts: How Language Models Merge Generated and Retrieved Contexts for Open-Domain QA?

### v2

@@ Abstract
While auxiliary information has become a key to enhance Large Language Models (LLMs), relatively little is known about how LLMs merge these contexts, specifically generated and retrieved. To study this, we formulate a systematic framework to identify whether LLMs’ responses, derived from the integration of generated and retrieved contexts, are attributed to either generated or retrieved contexts. To achieve this, we construct datasets with conflicting contexts, where each question is paired with both generated and retrieved contexts, yet only one of them contains the correct answer. Our experiments reveal a significant bias in LLMs (GPT-4/3.5 and Llama2) towards generated contexts, even when they provide incorrect information. We further identify two key factors contributing to this bias: i) contexts generated by LLMs typically show greater similarity to the questions, increasing their likelihood of selection; ii) the segmentation process used in retrieved contexts disrupts their completeness, thereby hindering their full utilization in LLMs. Our analysis enhances the understanding of how LLMs merge diverse contexts, offering valuable insights for advancing current augmentation methods for LLMs.
@@
Hexiang ${}^{\spadesuit\heartsuit}$, Fei ${}^{\spadesuit}$22footnotemark: 2  , Wanli ${}^{\spadesuit\diamondsuit}$, Yuanzhuo ${}^{\spadesuit}$, Qi ${}^{\spadesuit}$, Xueqi ${}^{\spadesuit\heartsuit}$$^{\spadesuit}$ Key Laboratory of AI Safety & Security, Institute of Computing Technology, Chinese Academy of Sciences, Beijing, China ${}^{\heartsuit}$niversity of Chinese Academy of Sciences, Beijing, China ${}^{\diamondsuit}$ankai University, Tianjin, China {tanhexiang21s, sunfei, wangyuanzhuo, caoqi, cxq}@ict.ac.cn
@@ 1 Introduction
Figure 1: Blue bars show the success number on NQ test set with only retrieved contexts, while warm orange bars depict the decline in success for the same questions when generated contexts are additionally incorporated.
@@
Building on the foundations laid by generation-augmented and retrieval-augmented methods, recent hybrid approaches have attempted to integrate them to further improve performance in tasks such as Question Answering (QA) Yu et al. (2022); Mallen et al. (2023). These hybrid approaches face a significant challenge: conflicts between diverse sources can impede the effectiveness of information integration Zhang et al. (2023). While recent works have investigated conflicts within contexts from a single source (Xie et al., 2023; Chen et al., 2022), it remains unclear how LLMs resolve conflicts between generated and retrieved contexts. This study, therefore, aims to investigate the underlying mechanisms by which LLMs process the two types of contexts, especially when they contain conflicting information.
@@
Our investigation was driven by a striking observation: in certain cases, models relying solely on retrieval contexts succeeded, whereas, counter-intuitively, hybrid approaches failed, as depicted in Figure 1. To uncover the underlying reasons, we proposed a systematic framework to dissect the process by which LLMs merge generated and retrieved contexts. We curated tailored datasets in which each question is accompanied by a pair of generated and retrieved contexts. These contexts are intentionally designed to be inconsistent with each other, and only one of them contains the correct answer to the corresponding question. These datasets provide an innovative lens to unravel how LLMs discern and integrate diverse contextual information in QA.
@@
In this paper, we conducted a series of controlled experiments using our uniquely designed datasets to empirically study this question, focusing on several state-of-the-art closed (GPT-3.5/4) and open (Llama2-7b/13b) LLMs. Surprisingly, our findings reveal a pronounced bias in LLMs to favor generated contexts, even when the generated contexts offer incorrect information while the retrieval contexts hold the correct answers. Furthermore, this bias persists regardless of whether the generated text was produced internally or by other LLMs. These findings highlight a critical challenge for existing LLMs in effectively merging internal parametric knowledge (i.e., generated contexts) and external information (i.e., retrieved contexts), under increasingly common non-tunable settings, e.g., those involving black-box APIs like GPT-4.
@@
To investigate the underlying causes of the bias, we study the role of confirmation bias Xie et al. (2023) (Section 5.1), text similarity (Section 5.2) and semantic completeness (Section 5.3). We find that: i) LLMs maintain a significant preference for generated contexts when they contain information inconsistent with LLMs’ parametric knowledge. ii) Generated contexts typically exhibit a higher degree of similarity to the questions compared to retrieved contexts, even when they contain incorrect information. This similarity gap is a significant factor in LLMs’ preference for generated contexts, which emphasizes the the need for caution with LLM-generated contexts, to avoid being misled by highly relevant but inaccurate information. iii) LLMs exhibit significant sensitivity to the semantic completeness of input contexts. The segmentation process used in retrieved contexts may disrupt their completeness, thereby hindering their full utilization in LLMs.
@@
This work preliminarily explores an growing challenge for LLMs to utilize contexts from diverse sources, especially as content generated by LLMs, which may include potential misinformation (Pan et al., 2023), becomes more prevalent on the internet. Furthermore, our findings offer valuable guidance for enhancing existing retrieval-augmented methods, such as optimizing passage segmentation in retrieval systems. Our main contributions can be summarized as:
@@
  • Our extensive analyses have identified two key factors, i.e., text similarity and semantic completeness, in the context utilization of LLMs. Moreover, the recently identified confirmation bias (Xie et al., 2023) plays a relatively minor role.
@@ 2 Background & Study Formulation
(a) Retrieval-Augmented Approach
(b) Generation-Augmented Approach
@@
Figure 2: The frameworks of retrieval-augmented approach, generation-augmented approach and hybrid approach.
@@ 2.1 Background
Retrieval-Augmented Approach. As shown in Figure 2, for a given question $q$ in a set of questions $\mathbb{Q}$, these approaches (Guu et al., 2020; Lewis et al., 2020; Ram et al., 2023; Gao et al., 2023) initially use a retrieval model $\gamma$ to select the top $k$ relevant documents $D_{k}^{\gamma}{=}\gamma_{k}(q,\mathbb{C}){=}\{d_{1}^{\gamma},\dots,d_{k}^{\gamma}\}$ from a corpus =$\mathbb{C}{=}\{d_{1},\dots,d_{|\mathbb{C}|}\}$. Then, a reader (often LLM) $\phi$ employs these documents $D^{\gamma}$ to generate an answer $a_{\phi}^{\gamma}$, expressed as $a_{\phi}^{\gamma}=\phi(q,D^{\gamma})$.
Generation-Augmented Approach. In contrast, as illustrated in Figure 2, these works (Yu et al., 2022; Sun et al., 2023; Liu et al., 2022) involve an LLM as a generator $\varrho$ to produce $k$ tailored background contexts $D_{k}^{\varrho}{=}\varrho_{k}(q){=}\{d_{1}^{\varrho},\dots,d_{k}^{\varrho}\}$ for a give question $q$, thereby enhancing the utilization of the LLM’s internal knowledge. These LLM-generated contexts $D^{\varrho}$ form the input for reader $\phi$ to produce the final answer: $a_{\phi}^{\varrho}=\phi(q,D^{\varrho})$.
@@ 2.2 Answer Tracing Task
Departing from previous research, our study investigates the mechanisms by which LLMs merge contexts from diverse sources, i.e., hybrid approach setting. As illustrated in Figure 3, we design a task to ascertain whether an answer $a_{\phi}$ originates from generated contexts $D_{k}^{\varrho}$ or retrieved contexts $D_{k}^{\gamma}$. Our task mirrors the basic process of hybrid approaches,while our focus is studying LLMs’ merging mechanisms by tracing the sources of the answers instead of improving QA performance. For a more controlled and simpler analysis, we limit the context to a single instance from each source, i.e., $k{=}1$ and $a_{\phi}{=}\phi(q,d^{\gamma}_{1},d_{1}^{\varrho})$. Subsequently, by comparing the answer $a_{\phi}$ with the answers derived from the retrieved context $a_{\phi}^{\gamma}$ and the generated context $a_{\phi}^{\varrho}$, we can determine its source, thereby analyzing the merging mechanism of LLMs.
@@ 3.1 Context-Conflicting Datasets
As depicted in Figure 4, in our dataset $\mathcal{D}_{\mathrm{cc}}$, each sample $x$ is a quintet $(q,d_{1}^{\gamma},d_{1}^{\varrho},a_{\phi}^{\gamma},a_{\phi}^{\varrho})$, where $d_{1}^{\gamma}$ is the context returned by retriever $\gamma$ for question $q$, $d_{1}^{\varrho}$ represents the context generated by LLM $\varrho$, $a_{\phi}^{\gamma}$ and $a_{\phi}^{\varrho}$ are the candidate answers provided by the reader $\phi$, each based solely on the respective contexts $d_{1}^{\gamma}$ and $d_{1}^{\varrho}$. To guarantee that our dataset is suitable for controlled experiments aimed at investigating the merging mechanisms of LLMs, it must adhere to specific criteria:
  • [itemsep=3pt, topsep=0pt, partopsep=0pt, parsep=2pt, leftmargin=10pt]
  • Traceability: $a_{\phi}^{\gamma}$ and $a_{\phi}^{\varrho}$ must be supported by their respective contexts, $d_{1}^{\gamma}$ and $d_{1}^{\varrho}$.
  • Exclusivity: Only one of the contexts, $d_{1}^{\gamma}$ or $d_{1}^{\varrho}$ provides the correct answer, i.e., either $a_{\phi}^{\gamma}$ or $a_{\phi}^{\varrho}$ matches the gold answer of question $q$.
@@
Context Preparation. Step 1 in Figure 4 illustrates the process of preparing retrieved and generated contexts for each question. For retrieved context, it is obtained from the top-$1$ ranked passage from Wikipedia using a retrieval model, e.g., Contriever (Izacard et al., 2021), a powerful off-the-shelf retrieval model that is extensively employed in various retrieval-augmented generation systems Shi et al. (2023); Ram et al. (2023).
For generated context, we follow the GenRead Yu et al. (2022), instructing the generator, i.e., LLM like GPT-4, to generate a background document based on the question. However, this method often yields contexts much longer (${>}250$ words) than the retrieved contexts (typically truncated to ${\sim}100$ words (Karpukhin et al., 2020; Izacard et al., 2021)). The discrepancy in length could potentially affect the merging mechanisms of LLMs, as discussed in Xie et al. (2023). To exclude this disturbance, we regulate the length of the generated context by incorporating length constraint in the prompt, resulting in an average length discrepancy below $3$%. All subsequent experiments, unless otherwise specified, employ this method to eliminate the impact of length variations. More details can be found in Appendix A.5.
@@
Building Context-Conflicting Dataset. Having obtained answers for each type of context, we are now positioned to construct our context-conflicting (CC) datasets, as depicted in Step 3 of Figure 4. Initially, We employ the exact match metric Yu et al. (2022) to evaluate the correctness of candidate answers derived from contexts, considering an answer correct if its normalized form matches any of the golden answers. Subsequently, the conflicting context datasets are composed of samples for which only one of the two types of contexts, either generated or retrieved, yields the correct answer, thereby ensuring the exclusivity. Notably, each dataset comprises two distinct subsets: AIG, consisting of samples with correct answers only in the generated context; and AIR comprising samples with correct answers only in the retrieved context.
@@ 3.2 Statistics of Datasets
For each reader-generator pair, we respectively constructs context-conflicting datasets from the original NQ and TQA test sets: NQ-CC (NQ-AIG + NQ-AIR) and TQA-CC (TQA-AIG + TQA-AIR).
We initially adopt a typical and simple setting in which a singular LLM serves as both the generator and reader. Table 1 provides detailed statistics for the constructed subsets corresponding to various LLMs, including GPT-4 (gpt-4-0613), GPT-3.5 (gpt-3.5-turbo-0613), Llama2-7b/13b (Llama-2-7b/13b-chat (Touvron et al., 2023)). The statistics show that the context-conflicting datasets form a substantial part of the test sets, underscoring the need to investigate how LLMs integrate these distinct contexts, as many questions are answerable with only one type. Notably, GPT-4 has fewer conflicting instances than other LLMs, because of its higher efficacy in answering questions using either solely retrieved or generated contexts.
@@ 3.3 Evaluation Metric
where $\mathtt{em}(a,b)$ returns 1 if $a$ exactly match $b$, and 0 otherwise. To facilitate a simple and efficient experiment, we define a synthesized metric as follows:
@@
The metric $\mathrm{DiffGR}$, ranging from $[-1,1]$, quantifies the extent of LLMs’ preference for choosing generated contexts over retrieved contexts. Using AIR as an example, where all correct answers come from retrieved contexts, an ideal $\mathrm{DiffGR}$ value should be $-1$, and $\mathrm{DiffGR}{>}0$ would indicate a significant bias towards generated contexts.
@@ 4 How LLMs Merge Contexts?
In this section, we conduct experiments on the proposed context-conflicting datasets to investigate the merging mechanism of the LLMs in hybrid approaches in different settings. We first consider a typical setting where generator and reader share a single LLM, to explore how LLMs merge retrieved and self-generated contexts (Section 4.2). Then, we extend our experiments to more flexible combinations of generator and reader (Section 4.2) to investigate their effects.

### v4

@@ Abstract
While auxiliary information has become a key to enhancing Large Language Models (LLMs), relatively little is known about how LLMs merge these contexts, specifically contexts generated by LLMs and those retrieved from external sources. To investigate this, we formulate a systematic framework to identify whether LLMs’ responses, derived from the integration of generated and retrieved contexts, are attributed to either generated or retrieved contexts. To easily trace the origin of the response, we construct datasets with conflicting contexts, i.e., each question is paired with both generated and retrieved contexts, yet only one of them contains the correct answer. Our experiments reveal a significant bias in several LLMs (GPT-4/3.5 and Llama2) to favor generated contexts, even when they provide incorrect information. We further identify two key factors contributing to this bias: i) contexts generated by LLMs typically show greater similarity to the questions, increasing their likelihood of being selected; ii) the segmentation process used in retrieved contexts disrupts their completeness, thereby hindering their full utilization in LLMs. Our analysis enhances the understanding of how LLMs merge diverse contexts, offering valuable insights for advancing current augmentation methods for LLMs [^1] Code released at https://anonymous.4open.science/r/7BB7/.
@@
Hexiang Tan♠♡, Fei Sun [^2] Corresponding author.  , Wanli Yang♠♢, Yuanzhuo Wang, Qi Cao, Xueqi Cheng♠♡ CAS Key Laboratory of AI Safety & Security, Institute of Computing Technology, Chinese Academy of Sciences, Beijing, China University of Chinese Academy of Sciences, Beijing, China Nankai University, Tianjin, China {tanhexiang21s, sunfei, wangyuanzhuo, caoqi, cxq}@ict.ac.cn
@@ 1 Introduction
Figure 1: Blue bars show the success number on the NQ test set with only retrieved contexts, while orange bars depict the decline in success for the same questions when generated contexts are additionally incorporated.
@@
Building on the foundations laid by generation-augmented and retrieval-augmented methods, recent hybrid approaches have attempted to integrate them to further improve performance in tasks like Question Answering (QA) Yu et al. (2022); Mallen et al. (2023). These hybrid approaches face a significant challenge: conflicts between diverse sources can impede the effectiveness of information integration Zhang et al. (2023). While recent works have investigated conflicts within contexts from a single source, either only retrieved (Chen et al., 2022) or generated (Xie et al., 2023), it remains unclear how LLMs resolve conflicts between generated and retrieved contexts. This study, therefore, aims to investigate the underlying mechanisms by which LLMs process the two types of contexts, especially when they contain conflicting information.
@@
Our investigation was driven by a striking observation: in certain cases, models relying solely on retrieval contexts succeeded, whereas, counter-intuitively, hybrid approaches failed, as depicted in Figure 1. To uncover the underlying reasons, we proposed a systematic framework to dissect the process by which LLMs merge generated and retrieved contexts. We curated tailored datasets in which each question is accompanied by a pair of generated and retrieved contexts. These contexts are deliberately designed to be inconsistent, with only one containing the correct answer to its corresponding question. These datasets provide a solid foundation for determining whether LLMs utilize retrieved or generated context to produce responses in QA tasks.
@@
In this paper, we conducted a series of controlled experiments using our uniquely designed datasets to empirically study this question, focusing on several state-of-the-art closed (GPT-3.5/4) and open (Llama2-7b/13b) LLMs. Surprisingly, our findings reveal a pronounced bias in LLMs to favor generated contexts, even when the generated contexts offer incorrect information while the retrieval contexts hold the correct answers. Furthermore, this bias persists regardless of whether the generated text was produced internally or by other LLMs (§4.2). These findings highlight a critical challenge for existing LLMs in effectively merging internal parametric knowledge (i.e., generated contexts) and external information (i.e., retrieved contexts), under increasingly common non-tunable settings, e.g., those involving black-box APIs like GPT-4.
@@
Furthermore, extensive controlled experiments are conducted to investigate the underlying causes of the bias and provide the following insights: (i) confirmation bias (Xie et al., 2023) is not a key factor5.1): LLMs maintain a significant preference for generated contexts when they contain information inconsistent with LLMs’ parametric knowledge. (ii) text similarity is a significant factor5.2): compared to retrieved contexts, generated contexts typically exhibit a higher degree of similarity to the questions, even when they contain incorrect information. The samples with a larger similarity gap between generated and retrieved contexts exhibit a more pronounced bias. These findings emphasize the need for caution with LLM-generated contexts, to avoid being misled by highly relevant but inaccurate information. (iii) semantic completeness matters5.3): LLMs tend to favor contexts with semantic integrity. The segmentation process used in retrieved contexts may disrupt their completeness, thereby hindering their full utilization in LLMs.
@@
This work preliminarily explores the growing challenge of LLMs utilizing contexts from diverse sources, especially in light of the increasing prevalence of LLM-generated content on the internet, which may contain potential misinformation (Pan et al., 2023). Furthermore, our findings offer valuable guidance for enhancing existing retrieval-augmented methods, such as optimizing passage segmentation in retrieval systems. Our main contributions can be summarized as:
@@
  • Our extensive analyses have identified two key factors, i.e., text similarity and semantic completeness, in the context utilization of LLMs. Moreover, we reveal that the confirmation bias (Xie et al., 2023) cannot account for the bias in this paper.
@@ 2 Background & Study Formulation
Figure 2: The frameworks of retrieval-augmented approach, generation-augmented approach, and hybrid approach.
@@ 2.1 Background
Retrieval-Augmented Approach. As shown in Figure 2, for a given question $q$ in a set of questions $\mathbb{Q}$, these approaches (Guu et al., 2020; Lewis et al., 2020; Ram et al., 2023; Gao et al., 2023) initially use a retrieval model $\gamma$ to select the top $k$ relevant documents $D_{k}^{\gamma}{=}\gamma_{k}(q,\mathbb{C}){=}\{d_{1}^{\gamma},\dots,d_{k}^{\gamma}\}$ from a corpus =$\mathbb{C}{=}\{d_{1},\dots,d_{|\mathbb{C}|}\}$. Then, a reader (often LLM) $\phi$ employs these documents $D^{\gamma}_{k}$ to generate an answer $a_{\phi}^{\gamma}$, expressed as $a_{\phi}^{\gamma}=\phi(q,D^{\gamma}_{k})$.
Generation-Augmented Approach. In contrast, as illustrated in Figure 2, these works (Yu et al., 2022; Sun et al., 2023; Liu et al., 2022) involve an LLM as a generator $\varrho$ to produce $k$ tailored background contexts $D_{k}^{\varrho}{=}\varrho_{k}(q){=}\{d_{1}^{\varrho},\dots,d_{k}^{\varrho}\}$ for a give question $q$, thereby enhancing the utilization of the LLM’s internal knowledge. These LLM-generated contexts $D^{\varrho}_{k}$ form the input for reader $\phi$ to produce the final answer: $a_{\phi}^{\varrho}=\phi(q,D^{\varrho}_{k})$.
@@ 2.2 Answer Tracing Task
Departing from previous research, our study investigates the mechanisms by which LLMs merge contexts from diverse sources in hybrid approaches. As illustrated in Figure 3, we design a task to ascertain whether an answer $a_{\phi}$ originates from generated contexts $D_{k}^{\varrho}$ or retrieved contexts $D_{k}^{\gamma}$. For a more controlled and simpler analysis, we limit the context to a single instance from each source, i.e., $k{=}1$ and $a_{\phi}{=}\phi(q,d^{\gamma}_{1},d_{1}^{\varrho})$. Then, by comparing the answer $a_{\phi}$ with the answers derived from the retrieved context $a_{\phi}^{\gamma}$ and the generated context $a_{\phi}^{\varrho}$, we can determine its source, thereby analyzing the merging mechanism of LLMs.
@@ 3.1 Context-Conflicting Datasets
As depicted in Figure 4, in our dataset $\mathcal{D}_{\mathrm{cc}}$, each sample $x$ is a quintet $(q,d_{1}^{\gamma},d_{1}^{\varrho},a_{\phi}^{\gamma},a_{\phi}^{\varrho})$, where $d_{1}^{\gamma}$ is the context returned by retriever $\gamma$ for question $q$, $d_{1}^{\varrho}$ represents the context generated by LLM $\varrho$, $a_{\phi}^{\gamma}$ and $a_{\phi}^{\varrho}$ are the candidate answers provided by the reader $\phi$, each based solely on the respective contexts $d_{1}^{\gamma}$ and $d_{1}^{\varrho}$. To guarantee that our dataset is suitable for controlled experiments aimed at investigating the merging mechanisms of LLMs, it should adhere to specific criteria:
  • Traceability: $a_{\phi}^{\gamma}$ and $a_{\phi}^{\varrho}$ should be supported by their corresponding contexts, $d_{1}^{\gamma}$ and $d_{1}^{\varrho}$.
  • Exclusivity: Only one of the contexts, $d_{1}^{\gamma}$ or $d_{1}^{\varrho}$, provides the correct answer, i.e., either $a_{\phi}^{\gamma}$ or $a_{\phi}^{\varrho}$ matches the gold answer of question $q$.
@@
Context Preparation. Figure 4 Step 1 illustrates the process of preparing contexts for each question. For retrieved contexts, it is obtained from the top-$1$ ranked passage from Wikipedia using Contriever (Izacard et al., 2021), a powerful off-the-shelf retrieval model that is extensively employed in various retrieval-augmented generation systems Shi et al. (2023); Ram et al. (2023).
For generated context, we follow the GenRead Yu et al. (2022), instructing the generator, e.g., LLM like GPT-4, to generate a background document based on the question. All LLMs in this paper, unless otherwise noted, have a temperature setting of zero to ensure result reproducibility. However, this method often yields contexts much longer (${>}250$ words) than the retrieved contexts (typically truncated to ${\sim}100$ words (Karpukhin et al., 2020; Izacard et al., 2021)). The discrepancy in length could potentially affect the merging mechanisms of LLMs (Xie et al., 2023). To exclude this disturbance, we regulate the length of the generated context by incorporating length constraint in the prompt, resulting in an average length discrepancy below $3$%. All subsequent experiments, unless otherwise specified, employ this method to eliminate the impact of length variations. More details can be found in Appendix A.2.
@@
Building Context-Conflicting Dataset. Having obtained answers for each type of context, we are now positioned to construct our context-conflicting (CC) datasets, as depicted in Step 3 of Figure 4. Initially, We employ the exact match metric Yu et al. (2022) to evaluate the correctness of candidate answers derived from contexts, considering an answer correct if its normalized form matches any of the golden answers. Subsequently, the context-conflicting datasets are composed of samples for which only one of the two types of contexts, either generated or retrieved, yields the correct answer, thereby ensuring the exclusivity. Notably, each dataset comprises two distinct subsets: AIG, consisting of samples with correct answers only in the generated context; and AIR comprising samples with correct answers only in the retrieved context.
@@ 3.2 Statistics of Datasets
For each reader-generator pair, we respectively construct context-conflicting datasets from test and dev sets of NQ and TQA: NQ-CC (NQ-AIG + NQ-AIR) and TQA-CC (TQA-AIG + TQA-AIR).
We initially adopt a typical and simple setting in which a singular LLM serves as both the generator and reader. Table 1 provides statistics for the constructed subsets corresponding to various LLMs, including GPT-4 (gpt-4-0613), GPT-3.5 (gpt-3.5-turbo-0613), Llama2-7b/13b (Llama-2-7b/13b-chat (Touvron et al., 2023)). The statistics show that the context-conflicting subsets form a substantial part of the datasets, underscoring the need to investigate how LLMs integrate these distinct contexts. Notably, GPT-4 has fewer conflicting instances than other LLMs, because of its higher efficacy in answering questions using either solely retrieved or generated contexts.
@@ 3.3 Evaluation Metric
where $\mathtt{em}(a,b)$ returns 1 if $a$ exactly match $b$, and 0 otherwise. The proportion of instances where $a_{\phi}$ does not match either $a_{\phi}^{\varrho}$ or $a_{\phi}^{\gamma}$ is negligible to the conclusion in this work, as demonstrated in Table 15. To facilitate a simple and efficient experiment, we define a synthesized metric as follows:
@@
The metric $\mathrm{DiffGR}$, ranging from $[-1,1]$, quantifies the extent of LLMs’ tendency to rely on generated contexts over retrieved contexts. Using AIR as an example, where all correct answers come from retrieved contexts, an ideal $\mathrm{DiffGR}$ value should be $-1$, i.e., LLMs should always rely on retrieved contexts.
@@ 4 How LLMs Merge Contexts?
This section conducts experiments on the developed datasets to investigate the merging mechanism of the LLMs in hybrid approaches in two settings. We first consider a typical setting where the generator and reader share a single LLM, to explore how LLMs merge retrieved and self-generated contexts (§4.1). Then, we extend our experiments to more flexible combinations of generator and reader (§4.2) to investigate their effects.